OpenStudy (anonymous):

find circumference. arc is 24 cm and angle is 135 degrees

OpenStudy (anonymous):

The EQUATION for calculating the ARC ANGLE RELATIONSHIP is theta/360° = s/C = s/2πr You have 135/360=24/C. Solving for C, you get 64 cm.